OpenStudy (surana):

Find the maximum or minimum of the following quadratic function: y = x^2 - 16

OpenStudy (surana):

\[x^{2}\]

OpenStudy (anonymous):

Is this a calculus question?

OpenStudy (anonymous):

if so you have to use first derivative test.

OpenStudy (surana):

Algebra II.

OpenStudy (anonymous):

graph it out.

OpenStudy (freckles):

The vertex can be found by writing your f(x)=x^2-16 in f(x)=(x-h)^2+k where this form tells us the vertex is (h,k)

OpenStudy (surana):

Okay.

OpenStudy (surana):

So what do I do?

OpenStudy (anonymous):